OpenStudy (anonymous):

A four sided, balanced die has the values 1, 2, 3, and 4. There are three successive throws. Find the probability of throwing 1, 2, and 4 in any order

OpenStudy (mathmate):

Hints: First throw: 3 choices (among 1,2 and 4) Second throw: 2 choices remaining Third throw: 1 choice remaining Success: 3*2*1 possibilities \(\Omega\): 4^3 possible outcomes.

OpenStudy (anonymous):

So would you do 3/4 * 2/3 * 1/2? Or would you divide 6/64?

OpenStudy (anonymous):

So would it be 3/4 * 2/4 * 1/4 since they are independent?

OpenStudy (perl):

$$ 111\\ 112\\ 113\\ 114\\ 121\\ 122\\ 123\\ 124\\ 131\\ 132\\ 133\\ 134\\ ... $$

OpenStudy (perl):

so we know there are 4*4*4 possible ways to list three integers from 1 to 4, since repetitions are allowed. Thus the denominator will be 4^3. For the numerator we have 3 ways to choose the first number, 2 ways for the second number, and 1 way for the last number. $$ \large \frac{3}{4} \cdot \frac{2}{4} \cdot \frac{1}{4}$$

OpenStudy (anonymous):

Ahhh okay I understand.

OpenStudy (perl):

so you can separate the problem into two parts. The numerator is counting how many ways you can rearrange 1,2,4 in any order 1,2,4 1,4,2 2,1,4 2,4,1 4,1,2 4,2,1

OpenStudy (perl):

the denominator is how many ways you can count 1,1,1 1,1,2 1,1,3 , etc

OpenStudy (anonymous):

So would the answer be 9.375%?